OpenStudy (anonymous):

please help if f(x)=(1/3)^x and a

OpenStudy (anonymous):

..are there any answer choices?

OpenStudy (anonymous):

Yes @jacquelyn1237 A.f(a)+ f(b)=3 B f(a) + 1/3=f(b) C f(a)= f(b) D. f (a) < f(b) E. f (a) >f(b)

Vocaloid (vocaloid):

first step is to narrow it down to D and E, since those are the only ones that make remote sense

Vocaloid (vocaloid):

an easy way to distinguish whether f(a) or f(b) is bigger is just to pick two numbers for a and b and see which one ends up bigger

Vocaloid (vocaloid):

if we let a = 1 and b = 2, see whether f(a) or f(b) is bigger.

OpenStudy (anonymous):

(you can rule out choices A and B because the questions asks which one "must" be true" and those only work for certain numbers)

OpenStudy (anonymous):

@Vocaloid which would be the a or the c in this?

Vocaloid (vocaloid):

?

Vocaloid (vocaloid):

just calculate f(a) when a = 1

OpenStudy (anonymous):

is the answer E, because its a whole fraccion?

OpenStudy (anonymous):

and how you know a its 1 ?

Vocaloid (vocaloid):

I just picked a value for a

Vocaloid (vocaloid):

the question just says a < b so I just picked a = 1 and b = 2 as test values

Vocaloid (vocaloid):

you can pick any two numbers for a and b as long as a < b

Vocaloid (vocaloid):

the point of me picking a = 1 and b = 2, was just to see whether f(a) or f(b) was bigger

OpenStudy (anonymous):

b is the exponent right?

Vocaloid (vocaloid):

f(x) = (1/3)^x f(a) = (1/3)^a f(b) = (1/3)^b find f(a) and f(b) if a = 1 and b = 2

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 3 } \right)^2=\frac{ 1 }{ 9 },\] \[\left( \frac{ 1 }{ 3 } \right)^3=\frac{ 1 }{ 27 }\] 2<3 \[\frac{ 1 }{ 9 } ?\frac{ 1 }{ 27 }\]

OpenStudy (anonymous):

yeah its true @surjithayer 1/9 is bigger than 1/27